OpenStudy (anonymous):

The mixture of glue and vinegar is 20kg and 30% of it is glue. It got added 22kg of vinegar afterwards. To make it 60% of vinegar, how much glue should we add?

hero (hero):

g = amount of glue in kilograms v = amount of vinegar in kilograms g + v = 20 g = .30(20) = 6 v = 14 I was able to determine this much about the problem. The part where it says "It got added 22kg of vinegar afterwards" That's where it starts getting fuzzy. Something about that sentence makes it seem like it was written by a fifth grader.

OpenStudy (anonymous):

well, its a fifth grader problem. and I translated it from another language so it will be hard to hard to understand, sorry

hero (hero):

Okay, I'll try my best.

OpenStudy (anonymous):

mmm, okay

hero (hero):

Seems like 22 kg of vinegar was added so: 6 + (14 + 22) = 42 6 + (36) = 42

hero (hero):

36 is the total amount of vinegar we have now.

OpenStudy (anonymous):

yeah, ok

hero (hero):

We only have 6 kg of glue

OpenStudy (anonymous):

right

hero (hero):

So now we need to figure out how much glue we need to add for the mixture to be 60% vinegar.

OpenStudy (anonymous):

yes, and that's where I got stuck

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok, alright

hero (hero):

But wait...something isn't adding up here.

OpenStudy (anonymous):

Ok, what?

hero (hero):

The vinegar is going to be 60% of the total kg amount

OpenStudy (anonymous):

yeah, that's correct.

hero (hero):

But something still doesn't make sense about it.

hero (hero):

I'm still trying to figure it out

OpenStudy (anonymous):

I know! I think its a writing error.

hero (hero):

No, unfortunately it's not.

OpenStudy (anonymous):

So, it's that hard.

hero (hero):

No, it's not. We're both just making it hard.

OpenStudy (anonymous):

How?

hero (hero):

I think we have to solve it by setting up a ratio

OpenStudy (anonymous):

Ok, let's try.

hero (hero):

The mixture of glue and vinegar is 20kg and 30% of it is glue. It got added 22kg of vinegar afterwards. To make it 60% of vinegar, how much glue should we add? Originally The ratio of glue to vinegar is \[\frac{6}{14}\] So 30% of the mixture is glue 70% is vinegar. After adding 22 kg of vinegar we have \[\frac{6}{36}\]

OpenStudy (anonymous):

yeah, so, is that the last answer?

hero (hero):

Now we have to add a certain amount of glue for the percentage to be 40% glue and 60% vinegar \[\frac{6 + g}{36} = \frac{40%}{60%}\] @babybluepies , patience...We're figuring this out together.

hero (hero):

\[\frac{6 + g}{36} = \frac{.40}{.60}\]

hero (hero):

That should be the right setup. I had that the first time but my percentage symbols didn't show up

OpenStudy (anonymous):

Thanks! i think I figured it out now, you really are a hero!

hero (hero):

You should be able to use that to solve for g

hero (hero):

Let me know what you get for g

OpenStudy (anonymous):

Okay, give me a moment.

OpenStudy (anonymous):

Kay, I got the answer but the thing is its a negetive number is that okay? The answer is 4.

OpenStudy (anonymous):

-4

hero (hero):

No, there's no possible way to get that. You have some issues with your algebra.

OpenStudy (anonymous):

Okay, harsh,but I'll check it again.

hero (hero):

All you had to do is cross multiply and simplify. Then isolate g.

hero (hero):

I did all the hard work for you already.

OpenStudy (anonymous):

Ok, sorry.

hero (hero):

I think you are stuck

hero (hero):

I'm going to finish solving it. I thought you had it from here.

hero (hero):

\[\frac{6 + g}{36} = \frac{.40}{.60}\] Cross multiply: .6(6 + g) = .4(36) 3.6 + .6g = 14.4 .6g = 14.4 - 3.6 .6g = 10.8 g = 10.8/.6 g = 108/6 g = 18